OpenStudy (anonymous):

Can someone please check my answers? What is the slope intercept form of the equation 3x-y= - 6? What is the slope intercept form of the equation x-3y=3? What is the slope intercept form of the equation 9x+3y=6? What is the slope intercept form of the equation x+3y=-6 A= y=3x+6 B. y=1/3x-1 C. y-3x+2 D. y+1/3x+2

OpenStudy (anonymous):

D. I mean y=1/3x-2

OpenStudy (anonymous):

@ranga are these right?

OpenStudy (ranga):

In C) the = sign is missing: y = -3x + 2 d) is incorrect. A and B are correct.

OpenStudy (anonymous):

C -1/3x-2

OpenStudy (ranga):

Yes, but it is D not C. D) y = -1/3x - 2

OpenStudy (anonymous):

okay so it sholud be A,D,C,B ?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Aren't A,B,C,D in the question YOUR answers?

OpenStudy (anonymous):

yes I was saying is that the order they go in

OpenStudy (ranga):

The slope intercept form of the equation: 3x - y = - 6 is y = 3x + 6 x - 3y = 3 is y = 1/3x - 1 9x + 3y = 6? is y = -3x + 2 x + 3y =-6 is y = -1/3x - 2

OpenStudy (anonymous):

okay yeah thats what I have, thank you for helping me.

OpenStudy (ranga):

sure.